INFRARED SPECTRA OF HYDROGEN - OCS CLUSTERS

\underline{JIAN TANG} AND A.R.W. McKELLAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

We have recently observed rich infrared spectra of helium clusters seeded with OCS, N_{2}O, or CO molecules. For He_{N}-OCS [1], transitions have been assigned up to N=8, for He_{N}-N_{2}O up to N=6, and for He_{N}-CO up to N=14. In all cases, lines due to even larger clusters are clearly present in the spectra but not yet definitively assigned. Using similar experimental conditions, we have also searched for analogous spectra using other clustering molecules such as hydrogen and neon. In the case of (H_{2})_{N}-OCS clusters, we begin with a good knowledge of the binary complex, H_{2}-OCS [2]. It is then possible to identify likely candidates for transitions due to the clusters with N = 2, 3, and 4. One limitation is that we are not sure what kind of spectrum to expect for these clusters. In the case of H_{2}-OCS, the spectrum due to ortho-H_{2} completely dominates when normal H_{2} is used in the jet expansion, and it is difficult to observe the (somewhat different) spectrum due to the para species even when using a para-enriched sample [2]. Since normal H_{2} was used in the present cluster experiments, it is similarly likely that ortho-H_{2} dominates, but we do not know what effect the angular momentum of each ground state ortho-H_{2} will have on the cluster spectra, especially when coupling between these J=1 hydrogens is present for N ≥ 2. To help to get around this problem, we have also recorded spectra of (HD)_{N}-OCS clusters, since in this case we can assume that the HD molecules will all relax to the J=0 ground state. We will present predictions of possible microwave transitions of hydrogen - OCS clusters, since such observations could provide positive confirmations of our assignments. Spectra showing Ne_{N}-CO transitions for N=2 and 3 will also be shown.

References

[1] J. Tang, Y. Xu, A.R.W. McKellar, and W. Jäger, Science, 297, 2030 (2002).

[2] J. Tang and A.R.W. McKellar, J. Chem. Phys. 116, 646 (2002).